# Trans-Glasso: A Transfer Learning Approach to Precision Matrix Estimation

### v1

@@ 6 Simulations
Figure 2: Simulation results for Model I. The default setting is $n_{0}=300$, $n_{\text{source}}=1000$ and $h=40$. In the first experiment, we increase $n_{0}$ while fixing $n_{\text{source}}$ and $h$. In the second experiment, we increase $n_{\text{source}}$ while fixing $n_{0}$ and $h$. In the third experiment, we increase both $n_{0}$ and $n_{\text{source}}$ while increasing $h$. More specifically, we let $n_{\text{source}}=3n_{0}$, and $n_{0}=70$ when $h=10$, $n_{0}=150$ when $h=20$, $n_{0}=300$ when $h=30$, $n_{0}=600$ when $h=40$ and $n_{0}=1200$ when $h=50$. In the fourth experiment, we fix both $n_{0}$ and $n_{\text{source}}$ while increasing $h$. Each dot represents the empirical mean across $30$ repetitions and the vertical bar represents $\text{Mean}\pm\frac{2}{\sqrt{30}}\times\text{Standard Error}$.
Figure 3: Simulation results for Model II. The default setting is $n_{0}=750$, $n_{\text{source}}=2000$ and $h=40$. In the first experiment, we increase $n_{0}$ while fixing $n_{\text{source}}$ and $h$. In the second experiment, we increase $n_{\text{source}}$ while fixing $n_{0}$ and $h$. In the third experiment, we increase both $n_{0}$ and $n_{\text{source}}$ while increasing $h$. More specifically, we let $n_{\text{source}}=3n_{0}$, and $n_{0}=100$ when $h=20$, $n_{0}=200$ when $h=30$, $n_{0}=300$ when $h=40$, $n_{0}=500$ when $h=50$, $n_{0}=800$ when $h=60$, $n_{0}=1000$ when $h=70$ and $n_{0}=1200$ when $h=80$. In the fourth experiment, we fix both $n_{0}$ and $n_{\text{source}}$ while increasing $h$. Each dot represents the empirical mean across $30$ repetitions and the vertical bar represents $\text{Mean}\pm\frac{2}{\sqrt{30}}\times\text{Standard Error}$.
Figure 4: Simulation results for Model III. The default setting is $n_{0}=150$, $n_{\text{source}}=1000$ and $h=40$. In the first experiment, we increase $n_{0}$ while fixing $n_{\text{source}}$ and $h$. In the second experiment, we increase $n_{\text{source}}$ while fixing $n_{0}$ and $h$. In the third experiment, we increase both $n_{0}$ and $n_{\text{source}}$ while increasing $h$. More specifically, we let $n_{\text{source}}=4n_{0}$, and $n_{0}=15$ when $h=20$, $n_{0}=30$ when $h=30$, $n_{0}=80$ when $h=40$, $n_{0}=300$ when $h=50$, and $n_{0}=1000$ when $h=60$. In the fourth experiment, we fix both $n_{0}$ and $n_{\text{source}}$ while increasing $h$. Each dot represents the empirical mean across $30$ repetitions and the vertical bar represents $\text{Mean}\pm\frac{2}{\sqrt{30}}\times\text{Standard Error}$.
Figure 5: Simulation results when the informative set $\mathcal{A}$ is unknown. We set $n_{0}=300$ and $n_{\text{source}}=1000$ for Model I; $n_{0}=750$ and $n_{\text{source}}=2000$ for Model II; and $n_{0}=300$ and $n_{\text{source}}=1000$ for Model III. Each dot represents the empirical mean across $30$ repetitions and the vertical bar represents $\text{Mean}\pm\frac{2}{\sqrt{30}}\times\text{Standard Error}$.
@@ 6.1 Data Generation Models
  • Model III: We generate the shared component from an Erdos–Renyi graph. Specifically, let $\tilde{\Omega}_{ii}=5$, $1\leq i\leq d$, and $U_{ij}\sim\text{Bernoulli}(0.02)$, $1\leq i<j\leq d$. If $U_{ij}=1$, let $\tilde{\Omega}_{ij}=\tilde{\Omega}_{ji}\sim\text{Unif}[-3,3]$; otherwise, set $\tilde{\Omega}_{ij}=\tilde{\Omega}_{ji}=0$. Let $\mathcal{S}_{\tilde{\Omega}}=\{(i,j)\in[d]\times[d]:\tilde{\Omega}_{ij}\neq 0\}$ be the support of $\tilde{\Omega}$. For given $h$ and $0\leq k\leq K$, uniformly choose $h$ entries ($h+1$ if $h$ is odd) from $[d]\times[d]\setminus\tilde{\Omega}$, denoted as $\mathcal{S}_{\Gamma^{(k)}}$, such that $(i,j)\in\mathcal{S}_{\Gamma^{(k)}}$ if and only if $(j,i)\in\mathcal{S}_{\Gamma^{(k)}}$. Let $\Gamma^{(k)\star}_{ij}=u_{ij}\mathds{1}\{(i,j)\in\mathcal{S}_{\Gamma^{(k)}}\}$, $1\leq i,j\leq d$, where $u_{ij}\sim\text{Unif}[-3,3]$. Finally, let $\Omega^{(k)}=\tilde{\Omega}+\Gamma^{(k)\star}+\sigma I_{d}$, where $\sigma$ ensures $\gamma_{\min}(\Omega^{(k)})\geq 0.1$ for $0\leq k\leq K$.
@@ 6.2 Experimental Design
We conduct four main experiments per model to investigate Trans-Glasso’s performance under various conditions:
@@
Each experiment is repeated 30 times to obtain reliable averages and standard errors.
@@ 6.3 Comparison Methods
  • Glasso-Target: Applies graphical lasso [Friedman et al., 2008] only to the target data.
  • Glasso-Pooled: Combines all target and source data, applying graphical lasso on the pooled dataset.
  • Trans-CLIME: A transfer learning approach for precision matrix estimation by Li et al. [2023a], which assumes a sparse divergence matrix across sources.
@@ 6.4 Results
The results for Models I – III are shown in Figure 24. Trans-Glasso generally outperforms baseline methods. In Experiment 3, Trans-Glasso shows consistency across all models, whereas Glasso-Pooled and Trans-CLIME do not. Trans-CLIME performs better with small $h$, but its performance deteriorates as $h$ increases. This is because a small increase in $h$ can significantly increase the sparsity of the divergence matrix defined in (2) when the covariance matrix is not sparse, as discussed in Appendix A. Therefore, when the precision matrix is sparse but the covariance matrix is not, Trans-Glasso is more reliable and robust.
@@ 6.5 Experiments with Unknown Informative Set
We perform simulation experiments with unknown $\mathcal{A}$, using the same three models. We divide $[K]$ into $[K]=\mathcal{A}\cup\mathcal{A}^{c}$. For $k\in\mathcal{A}$, we set the sparsity level $h$ to be small, and for $k\in\mathcal{A}^{c}$, $h$ to be large. Specifically, for Model I, $h=20$ for $k\in\mathcal{A}$ and $h=600$ for $k\in\mathcal{A}^{c}$; for Model II, $h=30$ for $k\in\mathcal{A}$ and $h=600$ for $k\in\mathcal{A}^{c}$; for Model III, $h=10$ for $k\in\mathcal{A}$ and $h=300$ for $k\in\mathcal{A}^{c}$. We implement the Trans-Glasso-CV algorithm (Section 4.3) and compare it with other methods. We vary $|\mathcal{A}|$ from 00 to $K$ to observe performance changes. Each experiment is repeated $30$ times with different random seeds.
@@ 7.1 Gene Networks Data for Brain Tissues
We apply Trans-Glasso to detect gene networks in different tissues using the Genotype-Tissue Expression (GTEx) data [^1] https://gtexportal.org/home/. Following Li et al. [2023a], we focus on genes related to central nervous system neuron differentiation (GO:0021953). We use the same 13 brain tissues as Li et al. [2023a], treating one as the target and the other 12 as sources. We only use 10 out of 13 tissues as targets, avoiding 3 due to small sample sizes. See Table 3 of the supplementary materials of Li et al. [2023a] for the complete list of tissues and Table 1 for the target tissues. We remove genes with missing values in these tissues and compare Trans-Glasso with baseline methods by computing cross-validation prediction error as defined in (16).
Figure 6 presents the final result, using Glasso-Target, Glasso-Pooled, and Trans-CLIME as baselines as in Section 6. To compare results across tissues, we report prediction errors relative to Glasso-Target. Figure 6 shows that Trans-Glasso performs best on most tissues. The relative prediction error of Trans-Glasso is always much smaller than $1$, indicating that it performs significantly better than using target data alone and is robust to negative transfer. In comparison, Glasso-Pooled and Trans-CLIME can perform worse or similar to Glasso-Target on some tissues.
@@ 7.2 Protein Networks Data for AML
We apply our method to a protein networks dataset for Acute Myeloid Leukemia (AML) subtypes. Understanding protein relationships in cells is crucial in cancer studies, and graphical models help build these networks. Following Peterson et al. [2015], we analyze protein levels for $213$ newly diagnosed AML patients, [^2] The dataset is provided as the supplement to Kornblau et al. [2009]. classified by the FAB system. Although protein interactions may differ across subtypes, common AML-related processes suggest shared connections. Thus, transfer learning can enhance subtype estimation using data from other subtypes. We focus on $18$ proteins involved in apoptosis and cell cycle regulation, studying four subtypes: M0 (17 subjects), M1 (34 subjects), M2 (68 subjects), and M4 (59 subjects) Peterson et al. [2015], Kanehisa et al. [2012].
For each subtype, after applying Trans-Glasso, we obtain the final estimated graph by choosing $20$ edges with the largest absolute values in the estimated precision matrix, resulting in graphs with similar edge numbers as in Peterson et al. [2015]. The final result is shown in Figure 7. Comparing with Peterson et al. [2015], many edges are discovered in both studies. However, our estimated graphs are more similar across subtypes. M0, M1, and M2 have the same structures, while M4 differs in two edges and has stronger connections between proteins in BAD families. This is supported by Tzifi et al. [2012], which observed higher expression levels of BAD family proteins in AML subtypes M4, M5, and M6 [Tzifi et al., 2012, Table 1], indicating more active interactions.
@@ 8 Conclusion
We introduce Trans-Glasso, a novel transfer learning approach to precision matrix estimation, which addresses the limitations of small target sample sizes by leveraging related source data. Trans-Glasso operates through a two-step process: an initial estimation using multi-task learning, followed by refinement via differential network estimation. This methodology achieves minimax optimality in a wide range of parameter regimes. Through extensive simulations, we demonstrate that Trans-Glasso consistently outperforms baseline methods, showcasing its robustness and adaptability, particularly in high-dimensional settings with limited target samples.
Future research directions include extending Trans-Glasso to estimate other graphical models, such as Gaussian copula [Liu et al., 2009, 2012a], transelliptical [Liu et al., 2012b], functional [Qiao et al., 2019, Tsai et al., 2024, Zhao et al., 2024], and Ising models [Kuang et al., 2017], and studying inferential methods within the transfer learning framework.
Appendix A Comparison with Li et al. [2023a]
In this section, we make a more detailed comparison of the similarity assumption made in Li et al. [2023a] and our Assumption 1. Let $\Upsilon^{(k)}=\left(\Omega^{(0)}-\Omega^{(k)}\right)\Sigma^{(k)}=-\Psi^{(k)}\Sigma^{(k)}$, or equivalently $\Psi^{(k)}=-\Upsilon^{(k)}\Omega^{(k)}$. Li et al. [2023a] assumes $\Upsilon^{(k)}$ is column-wise sparse in $L_{p}$-norm. Our Assumption 1 and the similarity assumption in Li et al. [2023a] do not imply each other and are generally incomparable. However, our Assumption 1 can be preferable in some applications. First, while the divergence matrix $\Upsilon^{(k)}$ is motivated by the KL divergence between Gaussian distributions, Assumption 1 is structural and applies to any distribution. Second, Assumption 1 is naturally interpreted in Gaussian graphical models, unlike the similarity assumption in Li et al. [2023a]. Finally, a technical advantage of our assumption is that, while sparsity in $\Psi^{(k)}$ does not generally imply sparsity in $\Upsilon^{(k)}$, or vice versa, further assumptions on $\Sigma^{(k)}$ or $\Omega^{(k)}$ can establish one or both directions. For example, note that
@@
If both $\Upsilon^{(k)}$ and $\Omega^{(k)}$ are sparse, then $\Psi^{(k)}$ is also sparse. Similarly, if $\Psi^{(k)}$ and $\Sigma^{(k)}$ are sparse, then $\Upsilon^{(k)}$ is sparse. The key difference is that $\Upsilon^{(k)}\implies\Psi^{(k)}$ relies on $\Omega^{(k)}$ being sparse, while $\Psi^{(k)}\implies\Upsilon^{(k)}$ relies on $\Sigma^{(k)}$ being sparse. In graphical models, sparsity assumptions on the precision matrix are more common than on the covariance matrix, making the sparsity assumptions on $\Psi^{(k)}$ weaker than those on $\Upsilon^{(k)}$.
Appendix B Preliminary Lemmas

### v2

@@ 6.1 Data Generation Models
  • Model III: We generate the shared component from an Erdos–Renyi graph. Specifically, let $\tilde{\Omega}_{ii}=5$, $1\leq i\leq d$, and $U_{ij}\sim\text{Bernoulli}(0.02)$, $1\leq i<j\leq d$. If $U_{ij}=1$, let $\tilde{\Omega}_{ij}=\tilde{\Omega}_{ji}\sim\text{Unif}[-3,3]$; otherwise, set $\tilde{\Omega}_{ij}=\tilde{\Omega}_{ji}=0$. Let $\mathcal{S}_{\tilde{\Omega}}=\{(i,j)\in[d]\times[d]:\tilde{\Omega}_{ij}\neq 0\}$ be the support of $\tilde{\Omega}$. For given $h$ and $0\leq k\leq K$, uniformly choose $h$ entries ($h+1$ if $h$ is odd) from $d]\times[d]\setminus\tilde{\Omega}$, denoted as $\mathcal{S}_{\Gamma^{(k)}}$, such that $(i,j)\in\mathcal{S}_{\Gamma^{(k)}}$ if and only if $(j,i)\in\mathcal{S}_{\Gamma^{(k)}}$. Let $\Gamma^{(k)\star}_{ij}=u_{ij}\mathds{1}\{(i,j)\in\mathcal{S}_{\Gamma^{(k)}}\}$, $1\leq i,j\leq d$, where $u_{ij}\sim\text{Unif}[-3,3]$. Finally, let $\Omega^{(k)}=\tilde{\Omega}+\Gamma^{(k)\star}+\sigma I_{d}$, where $\sigma$ ensures $\gamma_{\min}(\Omega^{(k)})\geq 0.1$ for $0\leq k\leq K$.
@@ 6.2 Experimental Design
Figure 2: Simulation results for Model I.
For each model, we conduct four main experiments to evaluate Trans-Glasso under various conditions, repeating each 30 times to ensure reliable averages and standard errors.
@@ 6.3 Comparison Methods
  • Glasso-Target: Applies graphical lasso (11) to target data only.
  • Glasso-Pooled: Pools target and source data, then applies graphical lasso.
  • Trans-CLIME: A transfer learning method by (22) that assumes a sparse divergence matrix across sources.
@@ 6.4 Results
Figure 3: Simulation results for Model II.
Figure 4: Simulation results for Model III.
The results for Models I – III are shown in Figure 24. Trans-Glasso generally outperforms baseline methods. In Experiment 3, Trans-Glasso shows consistency across all models, whereas Glasso-Pooled and Trans-CLIME do not. Trans-CLIME performs better with small $h$, but its performance deteriorates as $h$ increases. This is because a small increase in $h$ can significantly increase the sparsity of the divergence matrix when the covariance matrix is not sparse, as discussed in Appendix A. Therefore, when the precision matrix is sparse but the covariance matrix is not, Trans-Glasso is more reliable and robust.
@@ 6.5 Experiments with Unknown Informative Set
Figure 5: Simulation results when the informative set $\mathcal{A}$ is unknown.
We perform simulation experiments with unknown $\mathcal{A}$, using the same three models. We divide $[K]$ into $K]=\mathcal{A}\cup\mathcal{A}^{c}$. For $k\in\mathcal{A}$, we set the sparsity level $h$ to be small, and for $k\in\mathcal{A}^{c}$, $h$ to be large. Specifically, for Model I, $h=20$ for $k\in\mathcal{A}$ and $h=600$ for $k\in\mathcal{A}^{c}$; for Model II, $h=30$ for $k\in\mathcal{A}$ and $h=600$ for $k\in\mathcal{A}^{c}$; for Model III, $h=10$ for $k\in\mathcal{A}$ and $h=300$ for $k\in\mathcal{A}^{c}$. We implement the Trans-Glasso-CV algorithm (Section 4.3) and compare it with other methods. We vary $|\mathcal{A}|$ from 0 to $K$ to observe performance changes. Each experiment is repeated $30$ times with different random seeds.
@@ 7.1 Gene Networks Data for Brain Tissues
We apply Trans-Glasso to detect gene networks across tissues using the Genotype-Tissue Expression (GTEx) data  [^1] https://gtexportal.org/home/. Following (22), we focus on genes involved in central nervous system neuron differentiation (GO:0021953) across 13 brain tissues, treating one as the target and the remaining 12 as sources. To avoid small sample sizes, we exclude three tissues from the target set, using only 10. A complete list of tissues is in Table 3 of (22). We remove genes with missing values and compare Trans-Glasso to baselines via cross-validation prediction error (15).
Figure 6 presents the results, benchmarking Trans-Glasso against Glasso-Target, Glasso-Pooled, and Trans-CLIME, as in Section 6. Prediction errors are reported relative to Glasso-Target for cross-tissue comparison. As shown in Figure 6, Trans-Glasso consistently outperforms baselines, with a relative prediction error well below 1, demonstrating superior performance and robustness to negative transfer. In contrast, Glasso-Pooled and Trans-CLIME sometimes perform comparably or worse than Glasso-Target.
@@ 7.2 Protein Networks Data for AML
We apply our method to a protein network dataset for Acute Myeloid Leukemia (AML) subtypes. Understanding protein relationships is crucial in cancer studies, and graphical models aid in constructing these networks. Following (34), we analyze protein levels from 178 newly diagnosed AML patients, [^2] Dataset provided as a supplement to (16). classified by the FAB system. While protein interactions may vary across subtypes, shared AML-related processes suggest common connections, making transfer learning beneficial. We focus on 18 proteins involved in apoptosis and cell cycle regulation across four subtypes: M0 (17 subjects), M1 (34), M2 (68), and M4 (59) (34; 15).
For each subtype, we apply Trans-Glasso and extract the final estimated graph by selecting 20 edges with the largest absolute values in the estimated precision matrix, ensuring comparability with (34). The results in Figure 7 reveal substantial overlap with (34), though our graphs exhibit greater structural similarity across subtypes. M0, M1, and M2 share identical structures, while M4 differs by two edges and shows stronger BAD family protein connections. This aligns with (43), which reported higher BAD protein expression in AML subtypes M4, M5, and M6 (43, Table 1), suggesting more active interactions.
@@ 8 Conclusion
We introduce Trans-Glasso, a novel transfer learning approach for precision matrix estimation that mitigates small target sample limitations by leveraging related source data. Trans-Glasso follows a two-step process: initial estimation via multi-task learning, followed by refinement through differential network estimation. This method achieves minimax optimality across various parameter regimes. Extensive simulations show that Trans-Glasso consistently outperforms baselines, demonstrating robustness and adaptability, especially in high-dimensional settings with limited target samples.
Future directions include extending Trans-Glasso to other graphical models, such as Gaussian copula (26), transelliptical (25), functional (35; 41; 49), and Ising models (18), as well as developing inferential methods within the transfer learning framework.
Funding Acknowledgments
C.M. was partially supported by the National Science Foundation under grant DMS-2311127 and the CAREER Award DMS-2443867.
Appendix A Comparison of Similarity Assumptions
In this section, we provide a more detailed comparison between our Assumption 1 and the similarity assumption used in (4). Let
$\Upsilon^{(k)}=(\Omega^{(0)}-\Omega^{(k)})\Sigma^{(k)}=-\Psi^{(k)}\Sigma^{(k)},$
or equivalently, $\Psi^{(k)}=-\Upsilon^{(k)}\Omega^{(k)}$. Li et al. (4) assume that $\Upsilon^{(k)}$ is column-wise sparse in the $L_{p}$-norm. While our Assumption 1 and the assumption in (4) do not imply each other and are generally not directly comparable, our assumption may be more suitable in certain applications.
First, whereas $\Upsilon^{(k)}$ is motivated by controlling the KL divergence between Gaussian distributions, our Assumption 1 is purely structural and applicable to general distributions. Second, our assumption offers a natural interpretation in Gaussian graphical models, unlike the divergence-based similarity assumption in (4). Finally, a technical advantage of our framework lies in its flexibility when establishing sparsity relationships. While sparsity in $\Psi^{(k)}$ does not necessarily imply sparsity in $\Upsilon^{(k)}$, and vice versa, additional structural assumptions on $\Sigma^{(k)}$ or $\Omega^{(k)}$ can relate the two. Specifically, we have
@@ Appendix A Comparison of Similarity Assumptions
If both $\Upsilon^{(k)}$ and $\Omega^{(k)}$ are sparse, then $\Psi^{(k)}$ is also sparse. Conversely, if both $\Psi^{(k)}$ and $\Sigma^{(k)}$ are sparse, then $\Upsilon^{(k)}$ must be sparse. A key distinction is that sparsity in $\Upsilon^{(k)}$ implies sparsity in $\Psi^{(k)}$ only when $\Omega^{(k)}$ is sparse, whereas sparsity in $\Psi^{(k)}$ implies sparsity in $\Upsilon^{(k)}$ only when $\Sigma^{(k)}$ is sparse. In graphical models, sparsity assumptions on the precision matrix $\Omega^{(k)}$ are far more common than on the covariance matrix $\Sigma^{(k)}$, making sparsity in $\Psi^{(k)}$ generally a weaker and more practical assumption than sparsity in $\Upsilon^{(k)}$.
Although a comprehensive theoretical comparison between our Assumption 1 and the similarity assumption in (4) is challenging, we provide some empirical insights in the following discussion. We begin by introducing the divergence distance used in (4) to quantify the difference between $\Omega^{(0)}$ and $\{\Omega^{(k)}\}^{K}_{k=1}$, defined as
$h_{\mathcal{D}}\coloneqq\max_{1\leq k\leq K}\left\{\max_{1\leq j\leq d}\left\|\Upsilon^{(k)}_{j,\cdot}\right\|_{p}+\max_{1\leq j\leq d}\left\|\Upsilon^{(k)}_{\cdot,j}\right\|_{p}\right\},$
where $\Upsilon^{(k)}_{j,\cdot}$ denotes the $j$-th row of $\Upsilon^{(k)}$ and $\Upsilon^{(k)}_{\cdot,j}$ denotes its $j$-th column. Here, $\|\cdot\|_{p}$ is the $L_{p}$ vector norm, for which Li et al. (4) establish theoretical guarantees when $p\in[0,1]$.
Figure 8: Empirical analysis of how $h_{\mathcal{D}}$ increases with $h$ under the three models described in Section 6.1, for different choices of $p\in[0,1]$.
We empirically examine how $h_{\mathcal{D}}$ grows as $h$ increases under the three models discussed in Section 6.1, considering different choices of $p\in[0,1]$. Specifically, we set $d=100$ and $K=5$. For each value of $h$, we randomly generate 100 sets of target and source precision matrices $\{\Omega^{(k)}_{t}\}_{k=0}^{K}$ for $t=1,\dots,100$. For each $t$, we compute $\{\Upsilon^{(k)}_{t}\}^{K}_{k=1}$ and the corresponding divergence distance $h_{\mathcal{D},t}$, then estimate the empirical mean and standard error of $\{h_{\mathcal{D},1},\dots,h_{\mathcal{D},100}\}$, denoted as $\widebar{h}_{\mathcal{D}}$ and $\text{SE}_{h_{\mathcal{D}}}$, respectively. We visualize the error bars as $\widebar{h}_{\mathcal{D}}\pm\frac{2}{\sqrt{100}}\,\text{SE}_{h_{\mathcal{D}}}$. Additionally, we report the results for various values of $p$ to examine the effect of different $L_{p}$ norms.
The results are presented in Figure 8. We observe that the pattern of how $h_{\mathcal{D}}$ increases with $h$ depends on the choice of $p$, which defines the $L_{p}$ norm. Notably, when $p=0$, $h_{\mathcal{D}}$ directly quantifies the sparsity level of the divergence matrices. In this case, even for small values of $h$, the sparsity of the divergence matrices can be relatively high, aligning with our earlier discussion.
These observations suggest that there are scenarios where $h$ is small while $h_{\mathcal{D}}$ remains large. However, this does not imply that Trans-Glasso consistently outperforms Trans-CLIME. In fact, both algorithms achieve minimax optimality within certain parameter spaces. Our primary point is that Trans-Glasso has the potential to perform better than Trans-CLIME under specific conditions. Intuitively, when $h$ is small but $h_{\mathcal{D}}$ is large, Trans-Glasso is more likely to perform well; conversely, when $h_{\mathcal{D}}$ is small but $h$ is large, Trans-CLIME may be favored. In practice, the choice between these similarity assumptions should be guided by domain-specific knowledge. A more comprehensive theoretical and empirical comparison between Trans-Glasso and Trans-CLIME is an important direction for future research.
Figure 9: Simulation results for Model IV.
Appendix B More Simulation Details
In this section, we provide more simulation details.
For the simulation for Model I, the default setting is $n_{0}=300$, $n_{\text{source}}=1000$ and $h=40$. In the first experiment, we increase $n_{0}$ while fixing $n_{\text{source}}$ and $h$. In the second experiment, we increase $n_{\text{source}}$ while fixing $n_{0}$ and $h$. In the third experiment, we increase both $n_{0}$ and $n_{\text{source}}$ while increasing $h$. More specifically, we let $n_{\text{source}}=3n_{0}$, and $n_{0}=70$ when $h=10$, $n_{0}=150$ when $h=20$, $n_{0}=300$ when $h=30$, $n_{0}=600$ when $h=40$ and $n_{0}=1200$ when $h=50$. In the fourth experiment, we fix both $n_{0}$ and $n_{\text{source}}$ while increasing $h$.
For the simulation for Model II, the default setting is $n_{0}=750$, $n_{\text{source}}=2000$ and $h=40$. In the first experiment, we increase $n_{0}$ while fixing $n_{\text{source}}$ and $h$. In the second experiment, we increase $n_{\text{source}}$ while fixing $n_{0}$ and $h$. In the third experiment, we increase both $n_{0}$ and $n_{\text{source}}$ while increasing $h$. More specifically, we let $n_{\text{source}}=3n_{0}$, and $n_{0}=100$ when $h=20$, $n_{0}=200$ when $h=30$, $n_{0}=300$ when $h=40$, $n_{0}=500$ when $h=50$, $n_{0}=800$ when $h=60$, $n_{0}=1000$ when $h=70$ and $n_{0}=1200$ when $h=80$. In the fourth experiment, we fix both $n_{0}$ and $n_{\text{source}}$ while increasing $h$.
For the simulation for Model III, the default setting is $n_{0}=150$, $n_{\text{source}}=1000$ and $h=40$. In the first experiment, we increase $n_{0}$ while fixing $n_{\text{source}}$ and $h$. In the second experiment, we increase $n_{\text{source}}$ while fixing $n_{0}$ and $h$. In the third experiment, we increase both $n_{0}$ and $n_{\text{source}}$ while increasing $h$. More specifically, we let $n_{\text{source}}=4n_{0}$, and $n_{0}=15$ when $h=20$, $n_{0}=30$ when $h=30$, $n_{0}=80$ when $h=40$, $n_{0}=300$ when $h=50$, and $n_{0}=1000$ when $h=60$. In the fourth experiment, we fix both $n_{0}$ and $n_{\text{source}}$ while increasing $h$.
For the simulation results when the informative set $\mathcal{A}$ is unknown. We set $n_{0}=500$ and $n_{\text{source}}=1500$ for Model I; $n_{0}=750$ and $n_{\text{source}}=2500$ for Model II; and $n_{0}=100$ and $n_{\text{source}}=500$ for Model III.
In Figure 25, each dot represents the empirical mean across 3030 repetitions, and the vertical bar represents $\text{Mean}\pm\frac{2}{\sqrt{30}}\times\text{Standard Error}$.
Appendix C Simulation with Smaller Scales
To investigate the impact of the precision matrix scale on simulation results, we conduct additional experiments in this section. The data are generated using the same setup as Model I in Section 6.1, with a single modification: we set $\tilde{\Omega}_{ij}=2\times 0.6^{|i-j|}\mathds{1}(|i-j|\leq 1)$ for $1\leq i,j\leq d$, and draw $u_{ij}$ values from $\text{Unif}[-1,1]$. This adjustment reduces the overall scale of the precision matrix, making it more comparable to the setting examined in the simulation experiments of (4). We refer to this data generation setup as Model IV. The results are presented in Figure 9. As shown, the outcomes closely resemble those of Model I in Figure 2, suggesting that the performance of the various algorithms remains generally robust to changes in the scale of the precision matrix.
Appendix D Ablation Study of the Refinement Step
We empirically assess the contribution of the refinement step by evaluating the performance of the initial estimator $\widecheck{\Omega}^{(0)}$, which is the direct output of Trans-MT-Glasso. By comparing the performance of $\widecheck{\Omega}^{(0)}$ with that of the final estimator $\widehat{\Omega}^{(0)}$, we analyze the impact of the refinement step via differential network estimation. We adopt the same simulation settings as in Section 6 and evaluate empirical performance across all three data generation models.
The results are presented in Figures 1012. As shown, Trans-Glasso consistently outperforms Trans-MT-Glasso, underscoring the crucial role of the refinement step in enhancing estimation accuracy.
Figure 10: Simulation results of Trans-MT-Glasso for Model I.
Figure 11: Simulation results of Trans-MT-Glasso for Model II.
Figure 12: Simulation results of Trans-MT-Glasso for Model III.
Appendix E Comparative Analysis of Protein Networks Data Using Trans-CLIME
We apply the Trans-CLIME algorithm proposed by (4) to the protein network data analyzed in Section 7.2. Consistent with the methodology described in Section 7.2, we select the top 20 edges exhibiting the largest absolute values in the estimated precision matrices. The resulting graphs are presented in Figure 13.
Comparing these results to those in Figure 7, we observe that both Trans-Glasso and Trans-CLIME produce identical graphs for subtypes M0 and M2, suggesting a high degree of similarity between these two subtypes in terms of their protein interaction networks. However, the graphs obtained via Trans-CLIME exhibit fewer shared edges overall, indicating that this method tends to highlight greater differences among subtypes. In contrast, the graphs estimated by Trans-Glasso display greater similarity across different subtypes.
Figure 13: Protein networks for four AML subtypes obtained by Trans-CLIME. Red edges are shared by all subtypes, blue edges are not.
This difference in outcomes can be attributed to the distinct assumptions underlying each algorithm. Specifically, Trans-Glasso assumes that most entries in the precision matrices are shared across subtypes, leading to more similar graph structures. In contrast, Trans-CLIME relies on the assumption that the $L_{p}$-norm of the divergence matrix is small, which does not necessarily imply structural similarity. Determining which assumption is more biologically relevant to protein interactions requires deeper biological investigation, a question we leave for future work.
Appendix F Preliminary Lemmas